OpenStudy (petlovers123):

what is the volume of a carton that has a length of 2 1/4 feet, width of 1 3/5 feet, and height of 2 1/3 feet.

OpenStudy (petlovers123):

@raffle_snaffle

OpenStudy (petlovers123):

please help will medal

OpenStudy (isaidavila):

Convertirse the mixed number into normal fractions 9/4, 8/5 and 7/3 So the volume is the multiplication of the fractions

OpenStudy (petlovers123):

so i have to multiply all of them

OpenStudy (isaidavila):

9/4*8/5*7/3=??

OpenStudy (petlovers123):

is the answer 8 1/4

OpenStudy (petlovers123):

i have another question ca you help me

OpenStudy (isaidavila):

I get 8 2/5

OpenStudy (petlovers123):

so the answer is 8 2/5

OpenStudy (isaidavila):

Yes

OpenStudy (petlovers123):

ok thx so much ihve another question can you help me

OpenStudy (isaidavila):

Ok

OpenStudy (petlovers123):

Divide 2/3 ÷ 1/6 = ____.

OpenStudy (petlovers123):

I got 4

OpenStudy (isaidavila):

Yes you are right

OpenStudy (petlovers123):

oh ok thats all i put 4

OpenStudy (michele_laino):

hint: you have to multiply the first fraction by the inverse of the second fraction: \[\frac{2}{3} \times 6 = ...?\]

OpenStudy (isaidavila):

Ok no problem

OpenStudy (petlovers123):

Thanks

OpenStudy (michele_laino):

:)

OpenStudy (petlovers123):

im fanning both of you

OpenStudy (michele_laino):

thanks! :)

OpenStudy (petlovers123):

hey since your here can you help me with another question